OpenStudy (anonymous):

Given F(x)=1/(x+4) Find the average rate of change of F(x), on interval [10,10+h] you answer will be an expression involving h I got (10, 1/14) and (10+h, 1/((10+h)+4) i just need to know hot to simplify it more...I know that i need to do m=y^2-y^1/x^2-x^1

OpenStudy (anonymous):

just would love the steps for it plz:)

OpenStudy (anonymous):

Avg Rate=y2-y1/x2-x1 so =(1/14+h -1/14)/h

OpenStudy (anonymous):

=-1/14(14+h)

OpenStudy (anonymous):

i got it wrong, it was -1/(14h+196) not sure how they got that..

OpenStudy (anonymous):

o nvm...:P sorry